OpenStudy (anonymous):

The computer system access password is 5 digits. If the digits in the password are allowed to repeat, what is the probability that the hacker will guess correctly on the first try?

OpenStudy (anonymous):

20/100 1/5 20%

OpenStudy (elonasushchik):

Correct way to answer this would be: of ways for password = 10^5 so P[ guess correctly on 1st try ] = 1 / 10^5 = 1 / 100,000